# Causal Learning for Trustworthy Recommender Systems: A Survey

### v1

@@ 3.3 Counterfactual Reasoning based Methods
Moving beyond causal effect inference, counterfactual reasoning focuses on discovering effects of interventions in hypothetical scenarios to address the “what if” questions. Derived from this concept, two influential tools have been widely used in CTRS: Counterfactual Augmentation based RS (CARS) and Counterfactual Inference based RS (CIRS).
@@ 3.3.1 Counterfactual Augmentation based RS (CARS)
Counterfactual augmentation refers to the process of enriching data within counterfactual scenarios. This usually involves intervened synthetic data points that represent plausible alternative outcomes under different conditions. For example, Zhang et al. Zhang et al. (2021a) introduced a module for this purpose to address missing data issues. It generates positive and negative samples by replacing non-essential features and essential features, respectively. By incorporating counterfactual samples, these CARS methods can be better equipped to handle diverse and unseen biases, noises and attacks, leading to reliable recommendations in practice. However, there remains a need for effective strategies to assess and ensure the quality of counterfactual samples.
@@ 3.3.2 Counterfactual Inference based RS (CIRS)
As another influential tool, counterfactual inference is commonly used to evaluate and leverage causal relationships, aiming to answer the “what if” question. These questions vary across different problems. For instance, to handle the popularity bias in RS, Wei et al. Wei et al. (2021) investigated the question: “what the ranking score would be if the model only uses item properties”. By measuring the causal effects of two alternative situations, they identified that the key to mitigating popularity bias lies in reducing the direct effect of a counterfactual RS, which assigns ranking scores without considering the user-item match. Accordingly, an effective and compatible CIRS framework is devised for a wide range of existing models to alleviate the impacts of popularity bias.
@@ 4 Open Problems and Future Directions
In this section, we highlight several open problems and directions for future research.
Unresolved trustworthiness challenges. As shown in Table 1, there exist several trustworthiness challenges that are critical but remain unresolved by CTRS methods, such as population bias and poisoning attacks. Although we have presented some preliminary ideas for addressing these challenges through lens of causal learning, there is a clear need for more in-depth research to develop comprehensive solutions.
Beyond causal prior knowledge. Most existing CTRS methods He et al. (2023); Zheng et al. (2021) rely on handcrafted causal analysis and pre-defined causal models. This significantly limits their practical effectiveness, especially in scenarios with intricate causal relationships. In causal learning, causal discovery is a fundamental topic, where various advanced techniques have been researched in recent years, including SAT Hyttinen et al. (2014) and $\rm A^{*}$ based methods Lu et al. (2021). However, they have not received much attention in CTRS. Hence, there is a growing need to incorporate advanced causal discovery techniques into CTRS for enhanced practicality and applicability.
Causal-oriented trustworthy evaluations. There has been limited research on evaluating RS performance from causal learning perspectives Sato et al. (2019). Developing advanced causal-oriented evaluations helps to bridge the gap between causal theory and practical RS assessment. Moreover, existing trustworthiness evaluations primarily focus on assessing recommendation results, while ensuring trustworthiness across all stages of the RS lifecycle is crucial. Hence, multi-stage trustworthy evaluations are in great demand.
Towards more generalized situations. When delving into theoretical details of causal effect inference and counterfactual reasoning based methods, one can find that they essentially rely on several fundamental assumptions (e.g., consistency assumption) Pearl (2009). These assumptions, however, may not always hold in practical settings. Therefore, it is critical to develop and introduce advanced causal frameworks that offer enhanced generalization capabilities, thereby accommodating a wider range of real-world scenarios.
@@ 5 Conclusions
Recommender systems have become an integral part of our daily lives, playing a crucial role in shaping user experiences. Specifically, there has been an increasing focus on the trustworthiness of RS, where a variety of Causal-oriented Trustworthy RS (CTRS) methods emerge using the advancement of causal learning to address fairness, robustness and explainability challenges in RS. However, no systematic reviews are available. This survey presents an overview of CTRS methods. We not only identify major trustworthiness challenges at each stage of the RS lifecycle but also clarify the necessity of causal learning, providing valuable insights into the practical development and enhancement of CTRS. In addition, we have collected and organized the representative open-source CTRS algorithms and datasets on https://github.com/JinLi-i/CTRS_resource.

### v2

@@ 3.3 Counterfactual Reasoning based Methods
Moving beyond causal effect inference, counterfactual reasoning focuses on discovering the effects of interventions in hypothetical scenarios to address the “what if” questions. Derived from this concept, two influential tools have been widely used in CTRS: Counterfactual Augmentation based RS (CARS) and Counterfactual Inference based RS (CIRS).
@@ 3.3.1 Counterfactual Augmentation based RS (CARS)
Counterfactual augmentation refers to the process of enriching data within counterfactual scenarios. This usually involves intervened synthetic data points that represent plausible alternative outcomes under different conditions. For example, Zhang et al. Zhang et al. (2021a) introduced a module for this purpose to address missing data issues. It generates positive and negative samples by replacing non-essential features and essential features, respectively. By incorporating counterfactual samples, these CARS methods can be better equipped to handle diverse and unseen biases, noise, and attacks, leading to reliable recommendations in practice. However, there remains a need for effective strategies to assess and ensure the quality of counterfactual samples.
@@ 3.3.2 Counterfactual Inference based RS (CIRS)
Counterfactual inference is commonly used to evaluate causal relationships, aiming to answer specific “what if” questions. These questions vary across different problems. For instance, to handle the popularity bias in RS, Wei et al. Wei et al. (2021) investigated “what would the ranking score be if the model only uses item properties”. By analyzing the recommendation mechanism with counterfactual cases, they identified that the key to debiasing is reducing the direct effect of popularity $I_{\rm pop}$ on the outcome $Y$ and focusing on the total indirect effect Pearl (2009). However, CIRS methods often involve assumptions to estimate counterfactual results, which may not always hold in practice, potentially limiting their generalization ability.
@@ 4 Open Problems and Future Directions
Unresolved challenges in fairness and robustness. As shown in Table 1, several critical trustworthiness challenges, such as population bias and poisoning attacks, remain insufficiently addressed. These issues lead to unfair performance disparities and increased vulnerability to malicious attacks. Thus, there is a clear need for targeted research from a causal perspective to analyze and develop tailored solutions.
Beyond causal prior knowledge. Most existing CTRS He et al. (2023) rely on handcrafted causal analysis and predefined causal models, limiting their effectiveness in complex cases. While causal discovery has developed advanced techniques (e.g., $\rm A^{*}$) Lu et al. (2021) in causal learning, it remains underexplored in CTRS. Incorporating these methods is crucial for improved practicality and applicability.
Causality-oriented trustworthy evaluations. As discussed in Section 2.5, research on evaluating RS from causal learning perspectives Sato et al. (2019) is limited. Advancing causal-oriented evaluations can bridge the gap between causal theory and practical RS assessment. Moreover, ensuring trustworthiness across all stages is crucial, highlighting the necessity of multi-stage trustworthy evaluations.
Causality-oriented trustworthy generative RS. The integration of generative models (e.g., diffusion) into RS Li et al. (2025); Yang et al. (2023) introduces new trustworthiness challenges, including the risk of generating biased and uninterpretable recommendation content. Future research should develop causal generative frameworks that apply causal inference to intervene in the biased generation process and leverage counterfactual reasoning for improved interpretability.
Causality enhanced LLM-based RS. Large Language Models (LLMs) have been successfully introduced to RS Lin et al. (2024), enhancing reasoning capabilities but also raising concerns about inherited stereotypes and bias Kotek et al. (2023) in LLMs. Research directions include using causal inference in fine-tuning to reduce bias propagation.
@@ 5 Conclusions
Recommender systems are integral to daily life. Growing concerns over trustworthiness have led to Causality-oriented Trustworthy RS (CTRS), which leverages causal learning to enhance fairness, robustness, and explainability. However, there are no dedicated surveys devoted to this area. This survey presents a comprehensive overview of CTRS, identifying key trustworthiness challenges across the RS lifecycle and connecting them to the corresponding causal learning based solutions. A comprehensive summarization and comparison of CTRS methods have been provided. In addition, we compile representative open-source CTRS algorithms and datasets at https://github.com/JinLi-i/CTRS_resource.